# Large Model Strategic Thinking, Small Model Efficiency: Transferring Theory of Mind in Large Language Models

### v2

@@ 4.2 Out of Sample Results: Contexts
Figure 5: Results for LLaMa2-70b pre-trained, LLaMa2-7b pre-trained, and LLaMa2-7b fine-tuned grouped by game with out-of-sample contexts. We adopt the same conventions as in Fig.3. We keep the structure of the payoffs identical for games both within and out-of-sample. Within each game, and for each context, we observe that the propensity to cooperate with the fine-tuned model always lies in between that of the pre-trained model of the same size and the pre-trained model of the largest size.
Our results for the out-of-sample contexts paint an intricate and complex picture of the effects of fine-tuning. It is immediately clear by way of visual inspection of figure 5 that results are more "well-behaved" than the results for the in-sample contexts, with no trace of overcorrection nor exacerbation. The stark tendency to defect of the larger model when facing the "sports" context is, at a glance, arguably the most surprising result. The model’s consistency in defecting all the time, except when playing Prisoner’s Delight, suggests a hostile understanding of the social frame leading to adversarial behavior. Nevertheless, in line with previous results, LLaMa-70b does shift its behavior when a non-cooperative context is paired with an anti-dilemma.
Figure 6: Improvement of the fine-tuned model over the in-sample games coupled with out-of-sample contexts.
Despite these promising results, Figure 6 shows average improvement is much lower than in the within-sample case. Remarkably, the highest level of improvement is reached when the game is Prisoner’s Delight and the context is "roomsharing". This is possibly because LLaMa2 takes cues from both game structure and context, and in this scenario, the two of them work synergistically to point towards cooperation being the optimal behavior. This theory would also explain why the lowest rate of improvement belongs to the Stag Hunt game played under the "ventcap" scenario: the coordination required to achieve the payoff dominant equilibrium and the social framing of the interaction clash with one another, leading the model to pick the risk-dominant equilibrium.
@@
4.3 Out of Sample Results: Game and Context
@@
For our final robustness test, we analyze the behavior of the models under consideration in a completely different game structure, with a tailor-made context to accompany it. Experiments conducted on games with different payoffs under within-sample contexts can be found in SI C. In this section, we instead endeavor to understand if the fine-tuned model could generalize its learned behavior beyond the boundaries of two-player games, demonstrating a clearer understanding of social dilemmas and strategic interaction. To accomplish this, we use the Public Good Game (PGG), or the Prisoner’s Dilemma with $n$-players, in order to once again capture the tension between Nash equilibrium and Pareto optimality. Additionally, this allows us to account for scenarios in which choice is not binary, but rather continuous.
Model Mean St. Dev Obs.
LLaMa2-7b 5.49 0.86 300
LLaMa2-7b-FT 5.92 0.99 300
LLaMa2-70b 7 0 300
Table 1: Summary statistics from the Public Pool Game. Mean refers to the mean donation, st. dev. to the standard deviation of mean donation, and obs. is the count of how many instances of the game we run per model.
@@
Results are presented in Table 1. It is immediately evident that LLaMa2-70b exhibits more pro-social behavior and a greater willingness to donate, whereas the 7b pre-trained model displays comparatively selfish tendencies. Notably, the fine-tuned model demonstrates significantly higher levels of pro-social behavior than its pre-trained counterpart, a difference that is statistically significant across all conventional thresholds. This finding suggests that, despite the absence of specific training scenarios involving the game or context, the fine-tuned model retains an enhanced understanding of cooperation, resulting in increased donation levels. Although higher donations may not align with more sophisticated strategic thinking, the friendly interaction environment suggests that the elevated pro-sociality is best attributed to the context-aware intelligence inherent in LLaMa2-70b, a characteristic which has been effectively transferred to the fine-tuned model.
Additional visualizations for all our out-of-sample studies can be found in SI D.II.
@@ 5 Conclusion
Throughout the course of this research paper, we have investigated the effectiveness of generating datasets from large pre-trained LLMs for the purpose of instructing smaller models belonging to the same family in tasks that involve theory of mind and strategic thinking. Smaller, strategy-savvy, aligned linguistic models could offer a less costly alternative to consumers without access to powerful graphical processing units without compromising quality of output or performance.
Our research was therefore primarily motivated by our goal of assessing whether smaller models can be fine-tuned to obtain and retain not just the enhanced strategic capabilities of their larger siblings, but also their humanlike attention to social and contextual clues. Concurrently, our attempts to create a dataset to conduct this training find justification in the absence of a canonical corpus of texts that could be used for this purpose. Given LLaMa2-70b’s proven track record as an algorithm capable of balancing strategic considerations with social contexts and clues, we elected to use its responses to a set of 20 unique scenarios as training data. This training data was then used for fine-tuning a smaller algorithm, LLaMa2-7b, by using a memory-efficient technique known as LoRA. In our experiments, the small fine-tuned model departs significantly from the behavior of its pre-trained counterpart and imitates the behavior of a larger pre-trained model of the same family, maintaining considerably high levels of performance even on tasks and when facing contexts that were not present in the training sample.
The significance of these findings is threefold: First, it means that when given the right data in sufficient quantities, a small model can be tuned to mimic the human-aligned strategic behavior of a larger one. Second, it implies that LLMs can themselves be used to generate datasets for ToM tasks, or tasks for which a complete and canonical dataset does not exist. Third, and perhaps most significant, this suggests that fine-tuned models can efficiently and effectively learn to generalize from examples in the training data and display heightened performances even in out-of-sample scenarios and games.
Moreover, this research lays the groundwork for advanced agent-based modeling and simulation techniques, offering new insights into the emergence of pro-social behaviors, such as cooperation [53, 54], trust [55], and fairness [56] in complex systems. By demonstrating the potential of fine-tuned language models as strategic agents, our findings open avenues for creating computationally efficient digital twins of collective human behavior. This approach not only enhances our understanding of cooperation, trust, and fairness in social dynamics but also provides a powerful tool for modeling and analyzing intricate sociotechnical systems. Ultimately, this work bridges the gap between large language model capabilities and practical, scalable simulations of human and organizational behavior [57].
This study represents a promising step towards the development of smaller, aligned, and effective strategic language models. However, our approach does present some limitations. Firstly, the analysis was conducted using a single dataset of fixed size, leaving it unclear whether a larger or smaller dataset would yield different levels of effectiveness. Additionally, while we used a large representative sample of responses from all three models considered, the inherent volatility of high-temperature experiments suggests that a statistical evaluation of strategic choices would benefit from an even larger-scale experiment. Finally, it bears restating that our purpose has been to endow these models with the sophistication necessary to navigate the intersection of social dilemmas and social interactions, yielding results and outcomes that do not always align with what we would expect of a utility-maximizing player. For instance, when playing Prisoner’s Delight, some contexts were associated with an increased rate of defection post-fine-tuning, a choice inconsistent with traditional game theory. In light of this, researchers who would use fine-tuning to generate more sophisticated agents would do well to keep in mind the effect that context and framing inevitably have on these models’ decision-making processes.
In conclusion, this research highlights the need for further exploration into the intersection of language models, game theory, and social interactions. Whether LLMs will prove to be an asset for those wishing to study human behavior will also depend on these algorithms’ capabilities to conceptualize and react to extant structures of incentives which shape and inform interaction in our societies, and the extent to which fine-tuning can empower smaller models to do is yet to be fully understood. The development of small agile language models that can navigate complex social dilemmas has the potential to revolutionize fields such as behavioral economics, decision-making, and social science. As we continue to push the boundaries of what is possible with language models, we may uncover new and innovative ways to study and understand human behavior, and develop more effective and efficient algorithms that can be deployed in a wide range of relevant applications.

### v4

@@ 4.2 Out-of-Sample Results: Contexts
Figure 4: Improvement for the fine-tuned version of LLaMa2-7b on out-of-sample context and in-sample games grouped by (a) game or (b) context. We adopt the same conventions as in Fig.3. We keep the structure of the payoffs identical for games both within and out-of-sample. As opposed to within-sample scenarios, we observe no exacerbation or overcorrection.
Our exploration of out-of-sample contexts offers a detailed and nuanced understanding of the effects of fine-tuning, revealing both promising trends and inherent challenges. As illustrated in Figure 4 the outcomes in these contexts appear more "well-behaved" compared to the in-sample results, with no evidence of overcorrection or exacerbation. This suggests that fine-tuning contributes to greater stability when the model encounters new data. However, the average improvement observed in these out-of-sample contexts is noticeably lower than in the within-sample cases. This outcome is consistent with our expectations, reflecting a pattern commonly seen in traditional machine learning models where fine-tuned LLaMa2-7b demonstrates solid performance on novel data but falls short of perfection.
Notably, the most pronounced improvements are observed in scenarios featuring anti-dilemmas, such as the Prisoner’s Delight, and in contexts depicting interactions between friends. This finding implies that scenarios characterized by a higher degree of friendliness are more amenable to improvement, with "ventcap", the most antagonistic context, surprisingly ranking third in improvement. On the contrary, the severity of the dilemma itself does not seem to have a straightforward impact on the level of improvement. For instance, despite the Stag Hunt game offering stronger incentives for cooperation than the Prisoner’s Dilemma or Snowdrift, it ranks last in terms of improvement. This suggests a complex interplay between the nature of the dilemma and the effectiveness of fine-tuning, highlighting the need for a more granular understanding of how different contextual factors influence model performance.
@@
4.3 Out-of-Sample Results: Games
In order to assess whether the fine-tuned model has grasped the underlying dynamics of a certain game, we presented it with out-of-sample games and in-sample contexts. In this case, out-of-sample games are the exact games that the model has already seen in-sample, but with all payoffs scaled by a factor of 2. By scaling the payoffs rather than shifting them, we aim to capture a behavioral response to what is essentially a change in units of measure. Accordingly, we expect a strategically savvy model to keep its responses more or less constant across both in-sample and out-of-sample games. The results of our analysis can be found in Table 1, which reports our statistical tests of significance for observed differences. Notably, in most cases and on average, the fine-tuned model responds to the out-of-sample games in the same way it responded to the in-sample games. Of the 8 scenarios in which a statistically significant difference can be identified, 6 lie at the edge of significance, suggesting that the observed discrepancies might be attributed to noise or high temperature. We conclude that by keeping context fixed, the fine-tuned model can recognize game structures that are fundamentally identical to those already seen in training and act accordingly. This is evidence that fine-tuning can impart a higher level of sophistication on smaller models through the second-order effects of an improved Theory of Mind: the ability to recognize the dynamics underpinning a certain strategic interaction seamlessly extends to interactions that are fundamentally identical save for the unit of measure being used.
Scenario Normal C Ratio OoS C Ratio Difference SE $z$-score $p$-value
team_prison 0.75 0.71 $-0.04$ 0.02 1.6 $0.05^{*}$
team_delight 0.74 0.76 0.01 0.03 $-0.53$ 0.3
team_staghunt 0.74 0.71 $-0.03$ 0.03 1.18 0.12
team_snowdrift 0.70 0.72 0.02 0.03 $-0.88$ 0.19
IR_prison 0.74 0.71 $-0.03$ 0.03 1.05 0.15
IR_delight 0.75 0.70 $-0.05$ 0.02 2.0 $0.02^{*}$
IR_staghunt 0.72 0.69 $-0.03$ 0.03 1.29 0.10
IR_snowdrift 0.71 0.74 0.02 0.03 $-0.89$ 0.19
friendsharing_prison 0.75 0.70 $-0.05$ 0.02 2.0 $0.02^{*}$
friendsharing_delight 0.79 0.74 $-0.05$ 0.02 2.11 $0.02^{*}$
friendsharing_staghunt 0.71 0.77 0.06 0.03 $-2.17$ $0.01^{**}$
friendsharing_snowdrift 0.76 0.76 0.00 0.02 $-0.14$ 0.45
biz_prison 0.74 0.70 $-0.04$ 0.03 1.72 $0.04^{*}$
biz_delight 0.74 0.76 0.03 0.03 $-1.05$ 0.15
biz_staghunt 0.71 0.63 $-0.08$ 0.03 3.06 $0.00^{***}$
biz_snowdrift 0.74 0.76 0.01 0.03 $-0.53$ 0.30
environment_prison 0.63 0.66 0.03 0.03 $-0.96$ 0.17
environment_delight 0.69 0.72 0.03 0.03 $-1.0$ 0.16
environment_staghunt 0.65 0.64 $-0.02$ 0.03 0.61 0.27
environment_snowdrift 0.62 0.67 0.05 0.03 $-1.67$ $0.05^{*}$
AVERAGE 0.72 0.71 $-0.01$ 0.03 0.31 0.38
MEDIAN 0.74 0.71 $-0.01$ 0.03 1.05 0.15
Table 1: Difference in proportion $z$-score testing for propensity to cooperate in within-sample and out-of-sample games for the LLaMa2-7b fine-tuned model. For each scenario, we report: the proportion of cooperative choices in the within-sample game, the proportion of cooperative choices in the out-of-sample game, the difference in proportions, standard error, $z$-score, and associated $p$-value. Reported significance levels follow standard practices: one asterisk (*) for significance at the $0.05$ level, two asterisks (**) for significance at the $0.01$ level, and three asterisks (***) for significance at the $0.001$ level.
4.4 Out-of-Sample Results: Game and Context
Figure 5: Average contribution to the public good for each model, with standard errors in black.
@@ 4.4 Out-of-Sample Results: Game and Context
For our final robustness test, we analyze the behavior of the models under consideration in a completely different game structure, with a tailor-made context to accompany it. In this section, we endeavor to understand if the fine-tuned model could generalize its learned behavior beyond the boundaries of two-player games, demonstrating a clearer understanding of social dilemmas and strategic interaction. To accomplish this, we use the Public Good Game (PGG), or the Prisoner’s Dilemma with $n$-players, in order to once again capture the tension between Nash equilibrium and Pareto optimality. Additionally, this allows us to account for scenarios in which choice is not binary, but rather continuous.
@@
Results are presented in Figure 5. It is immediately evident that LLaMa2-70b exhibits more pro-social behavior and a greater willingness to donate, whereas the 7b pre-trained model displays comparatively more selfish tendencies. Notably, the fine-tuned model demonstrates increased levels of pro-social behavior compared to its pre-trained counterpart. This finding suggests that, despite the absence of specific training scenarios involving the game or context, the fine-tuned model retains an enhanced understanding of cooperation, resulting in increased contribution levels. Although higher donations may not align with more sophisticated strategic thinking, the friendly interaction environment suggests that the elevated pro-sociality is best attributed to the context-aware intelligence inherent in LLaMa2-70b, a characteristic that has been effectively transferred to the fine-tuned model. Additional visualizations for all our out-of-sample results can be found in the SI document, section C.II.
@@ 5 Conclusion
This research explores the potential of using large pre-trained language models (LLMs) to generate datasets for fine-tuning smaller models in tasks involving theory of mind (ToM) and strategic thinking. Our goal was to assess whether smaller, cost-efficient models could be trained to replicate not just the strategic capabilities of their larger counterparts, but also their nuanced understanding of social and contextual cues, offering a viable alternative to resource-intensive large models.
To address the absence of a canonical dataset for such tasks, we used responses from LLaMa2-70b to 20 unique scenarios as training data. This data was employed to fine-tune a smaller model, LLaMa2-7b, utilizing the state-of-the-art LoRA technique. Our experiments revealed that the fine-tuned smaller model could significantly deviate from its pre-trained behavior, closely imitating the strategic responses of the larger model. Remarkably, the fine-tuned model maintained high performance even in tasks and contexts not present in the training data.
The significance of our findings is threefold: (1) Smaller models, when provided with the right training data, can be fine-tuned to mirror the strategic and social behavior of larger models; (2) LLMs themselves can generate valuable datasets for tasks like ToM, where pre-existing datasets are unavailable; (3) Fine-tuned models exhibit strong generalization capabilities, performing well even in out-of-sample scenarios, which is crucial for their application in real-world settings.
This research paves the way for more efficient LLM-driven agent-based modeling and simulation techniques, which can be used in scenario studies to inform governance and policy decisions in complex sociotechnical systems. By demonstrating the potential of fine-tuned language models as strategic agents, we open up possibilities for creating computationally efficient digital twins of human collective behavior. This approach could significantly enhance our understanding of social dynamics and inform the design of more sophisticated models for studying human and organizational behavior.
Despite the promising results, the study has limitations. The fixed-size dataset used raises questions about the impact of dataset size on model effectiveness, and the variability of high-temperature experiments suggests that larger-scale statistical evaluations are warranted. Moreover, the model’s increased rate of defection in certain contexts, like Prisoner’s Delight, post-fine-tuning, highlights the complex influence of context and framing on decision-making processes. Future researchers should be mindful of these effects when developing sophisticated agents through fine-tuning.
In conclusion, this work underscores the potential of small, strategically aligned LLMs to navigate complex social dilemmas, offering valuable applications in fields like behavioral economics and social science. However, further research is needed to fully understand how fine-tuning can empower smaller models to engage with the incentive structures that drive human interactions. As we advance in this area, we may discover new methodologies for studying and simulating human behavior, leading to the development of efficient, context-aware algorithms for a wide range of applications.